OpenStudy (anonymous):

PLEASE HELP ME!! :'( I don't understand this. Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square. (5 points)

OpenStudy (amistre64):

you might wanna make it easier on yourself by creating some factors for it and expanding

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

x^2-6x+9 = 0 (x-3)^2 = 0 ----> x = 3 x^2-6x = -9 ----> (x-3)^2 - 9 = -9 ---> (x+3)^2 = 0 (6+-sqrt(6^2-4*9))/2 = 6/2 = 3 Not sure if this is what u need.

OpenStudy (anonymous):

That is PERFECT thank you so much. It show one of those things. thank you! :D

OpenStudy (anonymous):

No problem.